OpenStudy (anonymous):

Find the approximate solution of the given system of equations. 8x + 9y = 9 3x – 5y = 20 a. 3.36, –1.99 c. 0, –1.99 b. –1.11, 2.79 d. 3.36, 2.79

OpenStudy (anonymous):

8x+9y=9 3x-5y=20 The first step is to multiply the first equation by 5 and the second by 9 5(8x+9y=9) 9(3x-5y=20) Use the distributive property and then set them up like this and solve 40x+45y=45 + 27x-45y=180 --------------- 67x=225 67x/67=225/67 x= 3.36 Now plug x into either equation and solve. I chose the first as it is the easiest. 8(3.36)+9y=9 26.88 +9y=9 -26.88 -26.88 ------------------- 9y/9=17.88/9 y= -1.99 So the answer is A